# Variational Scarring in Graphene Quantum Dots

### v1

@@ Abstract
Most systems in nature are chaotic [1], and the conundrum of chaos in a quantum system has long been a topic of intense study [2, 3, 4, 5], dating back to the dawn of quantum mechanics [6, 7]. While the definition of quantum chaos is still evolving [8], it encompasses a diverse set of original phenomena, such as the scarring of the single-particle eigenstates [9, 10, 11]. In quantum scarring, the probability density of a quantum state is concentrated along a short, unstable periodic orbit (PO) of the corresponding chaotic classical system due to quantum interference, and the quantum state is imprinted by the corresponding PO. Since the initial prediction of quantum scars by Heller [9], extensive experimental attempts have been carried out to image these states. Even though analogues of scars have been observed in various classical wave experiments, such as in microwave resonators [12, 13], acoustic cavities [14], and fluid surface waves [15], their direct visualization in a genuine quantum system remained elusive until the recent success [16] in a stadium-shaped, graphene-based quantum dot (QD).
@@
Inspired by this progress, we here suggest an experimental pathway to directly visualize a different recently discovered member of the scar family, namely variational scars, first reported in disordered two-dimensional nanostructures [17]. Although these scars resemble the conventional scarring by Heller [9], they arise from a fundamentally different mechanism [17, 18]: These scarred eigenstates emerge as a synthesis of the underlying (near) degeneracy and the local nature of perturbation, guided by the variational principle. As a result, the scars in the perturbed quantum system form around the POs of the unperturbed classical counterpart, earning them the monikers perturbation-induced or variational. This scarring mechanism is ubiquitous [17, 18, 19, 20, 21], but experimental validation of variational scarring has yet to be achieved. To address this issue, we investigate the potential of a quantum dot on monolayer graphene to foster variational scars, thereby opening the door for a STM detection of these scarred states, in the same spirit as in Ref. [16].
@@
More specifically, we propose the following paradigmatic QD setup, schematically depicted in Fig. 1. The elliptical electrostatic confinement of the QD is achieved using a p-n junction within a heterostructure of monolayer graphene interfaced with sheets of hexagonal boron nitride that rests on a wafer composed of doped silicon covered with silicon oxide. A gate voltage $V_{g}$ can be applied to the underlying doped silicon substrate to tune the Fermi level $\varepsilon_{F}$ of the monolayer graphene. For imaging the shape of a quantum state, a grounded STM tip can be utilized, with a probing voltage bias $V_{s}$ applied via the graphene layer. Moreover, another tip can generate a local perturbation, creating a potential bump or dip depending on the applied voltage, within the QD confinement. This proposed QD device aligns not only with Ref. [16] but also with other recent advancements on quantum state control and imaging [22, 23]. We attribute the alluring opportunity of visualizing scars in the QD device to two pivotal developments: higher Q-value compared to foregoing experiments, such as the successfully imaging of a quantum corral [24], and an exquisite improvement in energy resolution up to the level of individual eigenstates, as shown in Refs. [22, 23].
@@
Figure 2: The figure summarizes the observed features of the variational scars in the elliptical QD perturbed by a nanotip. Subplot (a) displays a collection of high-energy eigenstates of the Hamiltonian in Eq. 1, with some strongly scarred by classical Lissajous periodic orbits. Subplot (b) highlights an exemplary scar, where the probability density concentrates along the Lissajous orbit in the shape of the symbol alpha (blue curve). The scar is also oriented in such a way that its loop avoids the bump (violet dot), effectively minimizing the perturbation caused by the nanotip. This scar originates from a relatively small subset of the eigenstates of the unperturbed system, as shown by the corresponding expansion in subplot (e). In addition, there are also variational scars that instead maximize the perturbation, like shown in subplot (d), and bouncing-ball states such as displayed in subplot (c). Subplot (f) illustrates the persistence of this type of scarring near to the optimal ellipticity of $\alpha=0.50$ against small deformations in QD confinement, when the underlying Lissajous orbits do not exist. Furthermore, subplot (g) demonstrates that the scars induced by the nanotip can endure under impurity noise, modeled as randomly located Gaussian bumps (green dots).
@@
$\begin{split}\mathcal{H}&=-t\sum_{\langle i,j\rangle}\left(a_{i}^{\dagger}b_{j}+b_{j}^{\dagger}a_{i}\right)\\ &+\sum_{i}\Big{[}V_{\textrm{ext}}+V_{\textrm{tip}}\Big{]}\left(a_{i}^{\dagger}a_{i}+b_{i}^{\dagger}b_{i}\right),\end{split}$ (1)
where $a$ ($a^{\dagger}$) and $b$ ($b^{\dagger}$) are the annihilation (creation) operators on the graphene sublattices A and B, respectively. We consider the conventional nearest-neighbor coupling with a hopping energy of $t=1.0$.
Regarding an effective description for the QD, previous experimental and theoretical studies have generally validated the usage of the harmonic approximation to represent the external confining potential of electrons in a QD [25], even extending to the quantum Hall regime under a strong magnetic field [26, 27] Therefore, the confinement potential is assumed to be parabolic, expressed as
@@
where $\omega_{0}$ is the overall confinement strength and $\alpha$ represents the ellipticity. We set $\omega_{0}=2$, aligning with reported experimental confinement strengths for similar QDs. Without loosing the generality, we focus an special value of $\alpha=0.5$ where accidental degeneracies occur in the unperturbed system associated with the Lissajous orbits existing in the classical counterpart. We note that Lissajous scars appear at various values of ellipticity $\alpha$ (see Ref. [18]). Moreover, as will be shown later, the ellipticity $\alpha$ does not need to precisely match the classical resonant conditions of the Lissajous orbits.
@@
$V_{\textrm{PI}}(\mathbf{r})=A\cdot\textrm{exp}\left[\frac{|\mathbf{r}-\mathbf{r}_{0}|^{2}}{\sigma^{2}}\right]$ (3)
@@
described by its amplitude $A$ and width $\sigma$. Similarly, randomly distributed Gaussian bumps are also employed to model impurity noise in a QD [31, 32, 33]. Here, we have selected the values of the amplitude and width so that the nanotip induces strong scarring in the studied energy regime. While we spotlight the local perturbation caused by a nanotip in this study, a similar effect can be engineered by introducing (an) adatom(s) at a specific location(s) on the graphene layer, or by leaving an n-doped region within the p-doped domain that forms the QD during the manufacturing phase. These alternative avenues elevate the tension caused by the feasibility of an arrangement of a STM tip with another nanotip might present. Nonetheless, we see a multi-tip approach as a viable possibility (see, e.g. Refs. [34, 35, 36]).
@@
Before diving deeper into the scarring occurring in the quantum dot, we want to address one more experimental aspect. In Fig. 1, we have also included the possibility of a homogeneous magnetic field which is perpendicular to the studied QD. The previous studies [18, 19, 20] that variational scars emerge in a disordered QD when the effective potential stemming from the external confinement and magnetic field satisfies the classical resonant condition for the Lissajous orbits. Alternatively, a variant of variational scarring can arise when spin-orbit coupling replaces local perturbation in a QD [37]. These systems also appear suitable for scar detection through STM [22, 23, 38, 39], and for identifying indirect scar signatures via NMR [40, 41] and spin-dependent transport [42, 43, 44, 45] measurements. However, we have opted to eliminate this kind of additional complexity and will concentrate on the scenario without the external magnetic field, which still offers the potential for direct visualizing variational scars, as demonstrated below.
@@
To compute thousands of eigenstates of our Hamiltonian, we employ the eigenfunction solver in the quantum transport software Kwant [46]. As expected based on the theory of varitional scarring [18, 20], when perturbed by the nanotip, multiple high-energy eigenstates of the elliptical QD show strong scarring, characterized by Lissajous orbits of the unperturbed classical system. In Fig. 2 (a), we show a collection of probability densities of the solved eigenstates, spotlighting an archetypal scar along with its classical PO and the bump in Fig. 2 (b). Moreover, we also observe quantum states, like shown in Fig. 2 (c) that feature classical “bouncing-ball-like” motion.
@@
To clarify, this type of scarred states do not exist in the unperturbed QD whose states $|\phi_{n}\rangle$ are instead expressible in terms of Hermitian polynomials.  [^1] The eigenstates of an elliptical are $|\phi_{n,m}\rangle\propto H_{m}(\sqrt{\alpha\omega_{0}}x)H_{n}(\sqrt{\omega_{0}}y)\exp[-\omega_{0}(\alpha^{2}x^{2}+y^{2})/2]$, where $H_{m}(\cdot)$ is the Hermite polynomial of order $m$. The corresponding energy spectrum shows degeneracies at commensurable frequencies, that is, when the ellipticity is a rational number. This observation strongly supports the assertion that the scars in the QD are variational in nature. To further confirm the origin of the reported scarring, we show two other trademarks unique to variational scarring. First, some of the scars align themselves with the bump, as shown in Fig. 2 (d), and thus maximize the perturbation caused by the nanotip; whereas the others, for instance, the scar in Fig. 2 (b) minimize it. Second, we expand the example scar in Fig. 2 (b) in terms of the eigenstates of the unperturbed elliptical QD as $|\textrm{Scar}\rangle=\sum_{n}c_{n}|\phi_{n}\rangle$. This type of “subspectrum” analysis of the expansion coefficients $|c_{n}|^{2}$ reveals that the scars are formed by a small set of (nearly) degenerate states. For instance, the showcased Lissajous scar is primarily ($\sim 90\,\%$) conjured by four eigenstates of the elliptical QD without the perturbation, as shown in Fig. 2 (e).
@@
Most importantly from the experimental point of view, the reported scars are robust. In Fig. 2 (f), we show that scars persist when the ellipticity $\alpha$ of the confinement is varied, counterintuively even under conditions where Lissajous orbits do not exist in the unperturbed classical system, a feature explained by the variational scarring theory [18] and discussed further below. Furthermore, even when the device is affected by noise, modeled as randomly scattered Gaussian bumps similar to nanotips but smaller in amplitude and width, some states remain distinctly scarred, as shown in Fig. 2 (e). However, the noise itself is able to trigger scarring [18], but here we focus on the scarring induced by the nanotip that provides an additional level of control over the generated scars.
@@
The focus on a single (main) perturbation also reveals another fascinating opportunity of variational scarring which has been passed over in the previous studies emphasizing multiple similar, randomly distributed impurities [18, 20, 19, 21, 17]. Generally, strong variational quantum scars are expected near commensurable frequency ratio $\omega_{x}/\omega_{y}=p/q$, denoted as $(p,q)$. Classical POs only exist when the frequencies are commensurable, meaning $\omega_{x}/\omega_{y}$ is a rational number, with $p$ and $q$ being coprime. Geometrically, the particle returns to its initial position with the original velocity after completing $p$ and $q$ oscillations between the vertical and horizontal turning points, respectively. Conversely, incommensurable frequencies lead to quasiperiodic motion, resulting in ergodic behavior on a torus [1].
@@
On the quantum side, (near) commensurable confinement frequencies give rise to special (nearly) degenerate states, referred as a “resonant” set, which are related a family of classical POs. Some linear combinations within this resonant set are scarred by these POs. When a moderate perturbation is applied, it generates eigenstates that are linear combinations of a single resonant set. According to the variational theorem, this process essentially involves extremizing the perturbation within the nearly degenerate subspace of the resonant set. Consequently, scarred states are favored because the perturbation is spatially localized. Therefore, the extremal eigenstates from each resonant set often exhibit scars corresponding to the associated POs.
@@
In contrast to the typical scenario involving numerous localized perturbations, we can selectively target scars of different geometries corresponding to distinct classical resonances based on the position of the nanotip. For instance, instead of the $(2,3)$ resonance scar shown in Figure 2(a) resembling the alpha shape, the ellipticity $\alpha=0.5$ would typically produce scars from the $(1,2)$ resonance, manifesting as either a parabola or a figure-eight. Both resonances are energetically close to each other. More explicitly, the $(1,2)$ resonance exhibits exact degeneracy, while the $(2,3)$ resonance is only approximately degenerate for $\alpha=0.5$. However, unlike the case of many random bumps [18], the scarred state from the $(2,3)$ resonance offers more effective extremization of the perturbation due to the precise placement of the nanotip. Notably, the self-crossing point of the $(2,3)$ resonance provides a particular advantage in minimizing and maximizing the perturbation (see, Figs. 2(a) and (d), respectively).
@@
Next, taking a step towards a first experimental visualization of the phenomenon, we demonstrate the opportunity for observing these scarred states in a STM measurement. For this purpose, we assess the local density of states (LDOS) at the probing bias $V_{s}$ as
$\rho(V_{s},\mathbf{r})=\frac{\eta}{\pi}\sum_{n}\frac{|\psi_{n}(\mathbf{r})|^{2}}{(\varepsilon_{F}+eV_{s}-\varepsilon_{n})^{2}+\eta^{2}},$ (4)
@@
employing the previously resolved eigenstates $\psi(\mathbf{r})$, and the energy resolution is set to be $\eta=0.0001$, in line with the previous experiments of the similar fashion [22, 23]. Moreover, the Fermi energy $\varepsilon_{F}$ is adjustable via the gate voltage $V_{g}$. The LDOS is directly linked to the differential conductance measured by a STM tip sweeping through a QD device.
@@
In particular, we focus on a situation where the probing bias $V_{s}$ is adjusted to match with the scarred state of Fig. 3. Figure 3 show the corresponding spatial distribution of the LDOS that paints a discernible pattern of resemble to the target scar. Our example underlies the point that the variational scars can indeed generate an observable STM signature. In other words, the STM can provide an experimentally feasible way to directly visualize the effect of variational scars, as opposed to the prevailing approach of deducing their presence incidentally through transport measurements, such as in Refs. [48, 49, 50, 51, 52].
@@
Figure 3: Figure shows the constructed LDOS map for a hypothetical differential conductance $\partial I/\partial V$ measurement. The probing bias $V_{s}$ of the simulated STM tip is adjusted to match with the example scar of Fig. 2 (b). The corresponding LDOS map contains a distinct imprint of the Lissajous orbit associated with the targeted scarred state.
@@
In addition to facilitating the observation of variational scars, the proposed approach may be also employed to explore the necessary but subtle consequence of quantum scars, namely antiscarring [53]. Moreover, an interesting avenue of future research is to investigate if there exist a hybrid form of variational [17, 20, 19, 18] and relativistic scars [54, 55, 56], or a possible transition between these two species of scarring as the energy is varied. Another interesting possibility is to utilize the presented quantum device as a platform to study quasi-scarred states in a spirit similar to that of Ref. [57] that might occur due to the roughness and atomic structure of the underlying monolayer graphene, and lattice scars within a more general framework in contrast to the previous study [58].

### v2

@@ Abstract
Most systems in nature are chaotic [1], and the conundrum of chaos in a quantum system has long been a topic of intense study [2, 3, 4, 5], dating back to the dawn of quantum mechanics [6, 7]. Although the subject of quantum chaos [8] has not yet reached its final form, it already encompasses a diverse set of original phenomena from the point of view of classical chaos, such as the scarring of the eigenstates of single particles [9, 10, 11].
In quantum scarring, the probability density of a quantum state is imprinted with a short, unstable periodic orbit (PO) of a chaotic classical system due to quantum interference. These scars are surprising as there is no such classical concentration of density near a PO in the long time limit for a bounded chaotic system, hence standing out against the corresponding, monotonous backdrop of quantum state randomness across most of phase space. Since the initial prediction of the phenomenon by Heller [9], extensive experimental attempts have been carried out to image scarred states. Even though analogues of scars have been observed in various classical wave experiments, such as in microwave resonators [12, 13], acoustic cavities [14], and fluid surface waves [15], their direct visualization in a genuine quantum system remained elusive until the recent breakthrough  [16] in a stadium-shaped, graphene-based quantum dot (QD).
@@
Inspired by this progress, we suggest an experimental pathway to directly visualize a different recently discovered member of the scar family, namely variational scars, first reported in disordered two-dimensional nanostructures [17]. Although these scars resemble the conventional scarring by Heller [9], they stem from a fundamentally different mechanism [17, 18]: the (near) symmetry of the unperturbed system leads to scarred states favored by the variational principle in the presence of a localized perturbation. As a result, the scars in the perturbed quantum system form around the POs of the unperturbed classical counterpart, so we call them perturbation-induced, variational scars. This scarring mechanism is ubiquitous [17, 18, 19, 20, 21, 22, 23, 24], but experimental validation of variational scarring has yet to be achieved. To address this issue, we investigate the potential of a realizable quantum dot on monolayer graphene to foster variational scars, allowing a STM detection of these scarred states, in the same spirit as in Ref. [16].
@@
We propose the following paradigmatic QD setup, schematically depicted in Fig. 1. The elliptical electrostatic confinement of the QD is achieved using a p-n junction within a heterostructure of monolayer graphene interfaced with sheets of hexagonal boron nitride that rests on a wafer composed of doped silicon covered with silicon oxide. A gate voltage $V_{g}$ can be applied to the underlying doped silicon substrate to tune the Fermi level $\varepsilon_{F}$ of the monolayer graphene. For imaging the shape of a quantum state, a grounded STM tip can be utilized, with a probing voltage bias $V_{s}$ applied via the graphene layer. Moreover, a second tip can generate a local perturbation, creating a potential bump or dip depending on the applied voltage, within the QD confinement. This proposed QD device aligns not only with Ref. [16] but also with other recent advancements on quantum state control and imaging [25, 26]. We attribute the alluring opportunity of visualizing scars in the QD device to two pivotal developments: higher Q-factor compared to foregoing experiments, such as the successfully imaging of a quantum corral [27], and an exquisite improvement in energy resolution up to the level of individual eigenstates, as shown in Refs. [25, 26].
@@
Figure 2: The figure summarizes the observed features of the variational scars in the elliptical QD perturbed by a nanotip. Subplot (a) displays a collection of high-energy eigenstates of the Hamiltonian described by Eq. 1, with some strongly scarred by classical Lissajous periodic orbits. Subplot (b) highlights an exemplary scar, where the probability density concentrates along the Lissajous orbit in the shape of the symbol alpha (blue curve). The scar is also oriented in such a way that its loop avoids the bump (violet dot), effectively minimizing the perturbation caused by the nanotip. These variational scars are accompanied by companion scars that, as shown in subplot (c), maximize the the effect of perturbation, rather than minimize it. Some of the displayed quantum states exhibit classical bouncing-ball-like motion (blue dashed circle), while others remain almost unaffected by the nanotip (red dashed circle).
@@
$\begin{split}\mathcal{H}&=-t\sum_{\langle i,j\rangle}\left(a_{i}^{\dagger}b_{j}+b_{j}^{\dagger}a_{i}\right)\\ &+\sum_{i}\Big{[}V_{\textrm{ext}}+V_{\textrm{tip}}+V_{\textrm{imp}}\Big{]}\left(a_{i}^{\dagger}a_{i}+b_{i}^{\dagger}b_{i}\right),\end{split}$ (1)
[…]
Previous experimental and theoretical studies have generally validated the use of the harmonic approximation to represent the external confining potential of electrons in a QD [29], even extending to the quantum Hall regime under a strong magnetic field [30, 31]. Therefore, the confinement potential is assumed to be an elliptical harmonic oscillator, expressed as
@@
where $\omega_{0}$ is the overall confinement strength and $\alpha$ represents the ellipticity. We set $\omega_{0}=2$, aligning with the experimental confinement strengths reported for similar QDs. Without loss of generality, we focus on a special value of $\alpha=0.5$ where, crucially, accidental degeneracies occur in the unperturbed system associated with the Lissajous orbits existing in the classical counterpart. We note that Lissajous scars appear at various values of ellipticity $\alpha$ (see Ref. [18]). Moreover, as will be shown later, the ellipticity $\alpha$ does not need to be precisely matched to the classical resonant conditions of the Lissajous POs.
@@
$V_{\textrm{tip}}(\mathbf{r})=A_{\textrm{tip}}\textrm{exp}\left[\frac{|\mathbf{r}-\mathbf{r}_{0}|^{2}}{\sigma_{\textrm{tip}}^{2}}\right]$ (3)
@@
described by its amplitude $A=0.35$ and width $\sigma=1.8$. Similarly, randomly distributed Gaussian bumps are also employed to model impurity noise in a QD [35, 36, 37]. Here, we have selected the values of the amplitude and width so that the nanotip induces strong scarring in the relevant energy regime. Although we spotlight the local perturbation caused by a nanotip in this study, a similar effect can be engineered by introducing an adatom(s) at a specific location(s) on the graphene layer, or by leaving an n-doped region within the p-doped domain that forms the QD during the manufacturing phase. These alternatives avoid the difficulty of arranging an STM tip in conjunction with the nanotip that generates the perturbation $V_{\textrm{tip}}$. Nonetheless, a multi-tip approach would provide a new level of controllability (see, e.g. Refs. [38, 39, 40]).
@@
In Fig. 1, we have included the possibility of a homogeneous magnetic field that is perpendicular to the QD. The previous studies, [18, 19, 20] variational scars emerged in a disordered QD when the effective potential stemming from the external confinement and magnetic field satisfies the classical resonant condition for the Lissajous orbits. Alternatively, a form of variational scarring can arise when spin-orbit coupling replaces local perturbation in a QD [23, 24]. These systems also appear suitable for scar detection through STM [25, 26, 41, 42], and for identifying indirect scar signatures via NMR [43, 44] and spin-dependent transport measurements [45, 46, 47, 48] . However, we will focus on the scenario without the external magnetic field, which still offers the potential for direct visualization of variational scars, as demonstrated below.
@@
We compute thousands of eigenstates of our Hamiltonian, utilizing the eigenfunction solver in Kwant [49]. As expected, based on the theory of varitional scarring, [18, 20], multiple high-energy eigenstates of the elliptical QD show strong scarring when perturbed by the nanotip, characterized by Lissajous orbits of the unperturbed classical system. In Fig. 2 (b) and (c), we show a collection of probability densities of the solved eigenstates, spotlighting two archetypal scars along with its classical PO and the bump in Fig. 2 (b) and (c). Moreover, the collection in Fig. 2 (a) includes quantum states with classical ”bouncing-ball-like” motion (blue dashed circle) as well as so-called remnants (green dashed circle), which appear to be almost unaffected by the applied perturbation.
@@
We emphasize that this type of scarred state does not exist in the unperturbed QD, whose eigenstates $|\phi_{n}\rangle$ are instead expressible in terms of Hermitian polynomials  [^2] The eigenstates of an elliptical are $|\phi_{n,m}\rangle\propto H_{m}(\sqrt{\alpha\omega_{0}}x)H_{n}(\sqrt{\omega_{0}}y)\exp[-\omega_{0}(\alpha^{2}x^{2}+y^{2})/2]$, where $H_{m}(\cdot)$ is the Hermite polynomial of order $m$. The corresponding energy spectrum shows degeneracies at commensurable frequencies, that is, when the ellipticity is a rational number.. This observation strongly supports the assertion that the scars in the QD are variational in nature. To further confirm the origin of the reported scarring, we show two other trademarks unique to variational scarring. First, some of the observed scars align themselves to minimize the perturbation caused by the nanotip, as seen in Fig. 2 (b); whereas the others counterintuivitely seek to maximize it, as displayed in Fig. 2 (c). Second, we expand the scar presented in Fig. 2 (b) in terms of the eigenstates of the unperturbed elliptical QD as $|\textrm{Scar}\rangle=\sum_{n}c_{n}|\phi_{n}\rangle$. This type of “subspectrum” analysis of the expansion coefficients $|c_{n}|^{2}$ reveals that the scars are formed by a small set of states close in energy, which is consistent with the theory of variational scarring [17, 20, 19, 18, 22]. For instance, the showcased Lissajous scar is primarily ($\sim 90\,\%$) conjured by three nearly-degenerate eigenstates of the elliptical QD without the perturbation, as shown in Fig. 3.
[…]
@@
Most importantly from the experimental point of view, the reported scars are robust. In Fig. 4 (a), we show that scars persist when the ellipticity $\alpha$ of the confinement is varied, counterintuitively even under conditions where the Lissajous POs do not exist in the unperturbed classical system, a feature explained by the variational scarring theory [18] and discussed further below. Furthermore, even when the device is affected by noise ($V_{\textrm{imp}}\neq 0$), modeled  [^3] More explicitly, the noise has form $V_{\textrm{imp}}(\bm{r})=A_{\textrm{imp}}\sum_{i}\exp\left(|\bm{r}-\bm{r}_{i}|^{2}/\sigma_{\textrm{imp}}^{2}\right)$ with $A_{\textrm{imp}}=0.15$ and $\sigma_{\textrm{imp}}=0.15$. In the energy range of interest, we have several bumps within the classically allowed area. as randomly scattered Gaussian bumps similar to the nanotip but smaller in amplitude (0.15) and width (0.8), some states remain distinctly scarred, as shown in Fig. 4 (b), which aligns with the behavior of bouncing-ball scars reported in Ref. [22]. This kind of noise is able to trigger scarring on its own [18], but here we focus on the scarring induced by the nanotip, providing an additional level of control over the generated scars.
Figure 4: Figure elucidates the robustness of the variational scars. Subplot (a) shows the persistence of this type of scarring near to the optimal ellipticity of $\alpha=0.50$ against small deformations in QD confinement, when the underlying Lissajous orbits do not exist. Furthermore, subplot (b) demonstrates that the scars induced by the nanotip (violet dot) can endure under impurity noise, modeled as randomly located Gaussian bumps (green dots). On the other hand, subplot (c) reveals the sensitivity of the scar associated with a specific Lissajous orbit to the location of the nanotip, as it strives to either maximize (upper row) or minimize (lower row) the perturbation. While the strength of the scarring depends on the position of the perturbation, the alpha-shape scars are present at the optimal ellipticity $\alpha=0.5$ across various tip positions.
@@
The focus here on a single (main) perturbation also reveals another fascinating opportunity of variational scarring which has been passed over in the previous studies emphasizing multiple similar, randomly distributed impurities [18, 20, 19, 21, 17]. Generally, strong variational quantum scars are expected near commensurable frequency ratio $\omega_{x}/\omega_{y}=p/q$, denoted as $(p,q)$. Classical POs only exist when the frequencies are commensurable, meaning $\omega_{x}/\omega_{y}$ is a rational number, with $p$ and $q$ being coprime. Geometrically, the particle returns to its initial position with the original velocity after completing $p$ and $q$ oscillations between the vertical and horizontal turning points, respectively. Conversely, incommensurable frequencies lead to quasiperiodic motion, resulting in ergodic behavior on a torus [1].
@@
On the quantum side, (near) commensurable confinement frequencies give rise to special (nearly) degenerate states, referred as a “resonant” set, which are related a family of classical POs. Some linear combinations within this resonant set are scarred by these POs. When a moderate perturbation is applied, it generates eigenstates that are linear combinations of a single resonant set. According to the variational theorem, this process essentially involves extremizing the perturbation within the nearly degenerate subspace of the resonant set. Consequently, scarred states are favored because the perturbation is spatially localized. Therefore, the extremal eigenstates from each resonant set often exhibit scars corresponding to the associated POs. The classical frequencies can be merely quasiresonant as long as the perturbation looms large enough to overcome the imperfect degeneracy.
@@
In contrast to the typical scenario involving multiple localized perturbations, we can selectively target scars of different geometries corresponding to distinct classical resonances based on the position of the nanotip. For instance, instead of the $(2,3)$ resonance scar shown in Figure 2(b) resembling the alpha shape, the ellipticity $\alpha=0.5$ would typically produce scars from the $(1,2)$ resonance, manifesting as either a parabola or a figure-eight. Both resonances are energetically close to each other. More explicitly, the $(1,2)$ resonance exhibits exact degeneracy, while the $(2,3)$ resonance is only approximately degenerate for $\alpha=0.5$. However, unlike the case of many random bumps [18], the scarred state from the $(2,3)$ resonance offers more effective extremization of the perturbation due to the precise placement of the nanotip. Notably, the self-crossing point of the $(2,3)$ resonance provides a particular advantage in minimizing and maximizing the perturbation (see, Fig. 2(b) and (c), respectively).
On the other hand, moving the nanotip affects the level of scarring, as illustrated in Fig. 4 (c), for three different tip positions shifted from the center of the QD by $\Delta x=2,6,10$ along the positive x-axis [the “optimized” scars shown in Fig. 2(b) and (c) have $\Delta x=10$]. The scars are weaker when the bump caused by the tip does not align with a position where the (1,2) resonance scar can optimally minimize or maximize the perturbation, i.e., by coinciding with or avoiding the self-crossing point of the scar and the bump. Nevertheless, noticeable concentrations of probability density along the corresponding Lissajous orbits are observed in some of the perturbed states. Furthermore, since the potential and, consequently, the periodic orbits scale with energy, the effect of the tip position can be mitigated by either focusing on a different energy range or slightly adjusting the confining potential.
@@
Next, taking a step towards a first experimental visualization of the phenomenon, we demonstrate the opportunity for observing these scarred states in STM measurements. For this purpose, we assess the local density of states (LDOS) at the probing bias $V_{s}$, which is directly linked to the differential conductance measured by a STM tip sweeping through a QD device, as
$\rho(V_{s},\mathbf{r})=\frac{\eta}{\pi}\sum_{n}\frac{|\psi_{n}(\mathbf{r})|^{2}}{(\varepsilon_{F}+eV_{s}+\Delta V-\varepsilon_{n})^{2}+\eta^{2}},$ (4)
@@
employing the previously resolved eigenstates $\psi(\mathbf{r})$, and the energy resolution is set to be $\eta$, in line with the previous experiments of the similar fashion [25, 26].
@@
In particular, we highlight a situation where the probing bias $V_{s}$ is adjusted to match with the scarred state of Fig. 2(b). Figure 5 (a) presents the corresponding spatial distribution of the LDOS that paints a discernible pattern resembling the target scar at moderate energy resolution $\eta=10^{-4}$ with $\Delta V=0$. Remarkably, the scar pattern remains apparent even when the probe voltage is mismatched, like in Fig. 5(b) with $\Delta V=3.5\cdot 10^{-4}$, or when the energy resolution is degraded to $\eta=10^{-3}$, as shown in Fig. 5(c). This simple example thus underlines the point that the variational scarring can indeed lead to a resilient observable STM signature. In other words, the STM can provide an experimentally feasible way to directly visualize the effect of variational scars, as opposed to the prevailing approach of deducing their presence incidentally through transport measurements, such as in Refs. [52, 53, 54, 55, 56].
@@
Figure 5: Figure shows the constructed LDOS map for a hypothetical differential conductance $\partial I/\partial V$ measurement. The probing bias $V_{s}$ of the simulated STM tip is adjusted to match with the example scar of Fig. 2 (b). As seen in (a), the corresponding LDOS map contains a distinct imprint of the Lissajous orbit associated with the targeted scarred state when $\Delta V=0$ and $\eta=10^{-4}$. The scar pattern persists even when $\Delta V=3.5\cdot 10^{-3}$ in (b) and $\eta=10^{-3}$ in (c).
@@
In addition to facilitating the observation of variational scars, the proposed approach may be also employed to explore a necessary but subtle consequence of quantum scarring, namely antiscarring [57, 58]. An interesting avenue of future research is to investigate if there exists a hybrid form of variational [17, 20, 19, 18] and relativistic scars [59, 60, 61], or a possible transition between these two types of scarring, as the energy is varied. Another interesting possibility is to utilize the presented quantum device as a platform to study quasi-scarred states in a spirit similar to that of Ref. [62] that might occur due to the roughness and atomic structure of the underlying monolayer graphene, and lattice scars within a more general framework in contrast to the previous study [63].